OpenStudy (anonymous):

A cable company charges a monthley fee of 24.50 for a cable service. There is an additional monthley fee of 3$ for each premium channel. George has a cable service with 4 premium channels. Write and evaluate a function to find the monthly cost of georges cable service

OpenStudy (anonymous):

@satellite73 :D I miss you

OpenStudy (anonymous):

if you put the number of premium channels as say \(x\) then the cost for the premium channels would be \(3x\) and the total cost would therefore be \(C(x)=24.50+3x\)

OpenStudy (anonymous):

i have been here, where have you been? http://www.youtube.com/watch?v=M80cMw15tzI

OpenStudy (anonymous):

haha same been there for a year now since we last talked ;)

OpenStudy (anonymous):

replace \(x\) by 4 and get \(C(4)=24.50+3\times 4\)

OpenStudy (anonymous):

i got a new car this year

OpenStudy (anonymous):

What?! No bike?!

OpenStudy (anonymous):

Wow! Nice!!

OpenStudy (anonymous):

I repainted my 69' camaro

OpenStudy (anonymous):

1966 signet (valiant) convertible what color?

OpenStudy (anonymous):

Dark blue with black stripes instead of white :) trying to be orignal

OpenStudy (anonymous):

Looks like a monster coming down the road

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (anonymous):

Mr. Parker is a car salesman. He makes a salary of 55,000 a year. He also earns a commision of 1/12 of his total sales for the year. This year, My Parker has a total of 295,000. write /evaluate

OpenStudy (anonymous):

\[P(x)=55,000+\frac{1}{12}x\] where as before \(x\) is the amount of sales then \[P(295,000=55,000+\frac{1}{12}\times 295,000\] whatever that is gotta run, good luck

OpenStudy (anonymous):

mr parker, lol